I need new and unique answers, please. (Use your own words, don't copy and paste), Please Use your keyboard (Don't use handwriting) Thank you..

Q1-

Abdulaziz Co. purchased a machine in 2013 for 50,000 that has a useful life of 5 years with a salvage value of 5,000.

Calculate the depreciation expense, accumulated depreciation, book value throughout its useful life using:

1- Straight-line Method.

2- Units of Production Method if the machine produces 100,000 units.

Here is a table of units produced each year:

First

Second

Third

Fourth

Fifth

23,000

25,000

-

30,000

22,000

3- Double Declining Balance Method (2 Marks).

Q2- On June 1, 2019, ABC Company signed a $25,000, 120-day, 6% note payable to cover a past due account payable.
a. What is the total amount of interest to be paid on this note?
b. Prepare ABC Company's general journal entry to record the issuance of the note payable
c. Prepare ABC Company's general journal entry to record the payment of the note on
September 29, 2019 (2 marks).

Q3. What are the characteristics of corporations (1 mark).
Answer #1

Straight line method

Depreciation per year= cost-salvage value/no of years of useful life

Depreciation per year=45000/5000=9000 per year

Accumulated depreciation and book value will be as follows

If u follow accumulated depreciation method book value will remain the same in the books with depreciation accumulated each year @9000 per year

Production units method

Depreciation amount should reflect the pattern in which income is accrued to the firm thus

Depreciation@1 year=45000*23000/100000=10350

Depreciation@2 year =45000*25000/100000=11250

Depreciation@3 year = 0

Depreciation@4 year =45000*30000/100000=12000

Depreciation@5 year=45000*22000/100000=8800

Interest will be 25000*6/100*120/360=375

Assuming 360 days a year

375 will be for 3 130 days period

Journal entry

Accounts payable a/c. Dr 25000

To 6% notes payable a/c. 25000

On the date of payment entry will

Notes payable a/c. Dr. 25000

Interest expense a/c Dr 375

To bank a/c. 25375